# From See to Shield: ML-Assisted Fine-Grained Access Control for Visual Data

### v1

@@ 6.1. PSO Detection Performance
Table 3(a) presents the performance of segmentation models, highlighting variations across object categories. SegFormer-B5 and DeepLabV3+ perform particularly well on categories including face, person, handwriting, and fingerprint, but fail in classes that have fewer samples in the training set, such as license plate, physical disability, and medical history. In contrast, YOLO-Seg, FPN-R, and FPN-X provide more balanced performance results in all classes, with FPN-X101 achieving the highest overall mIoU score. Our results indicate that all models exceed a certain threshold for face and person detection, reflecting knowledge already embedded in pre-trained weights and refined during fine-tuning. However, their performance drops significantly for classes such as physical disability, signature, and medicine, largely due to class imbalance. These categories are considered highly sensitive; therefore, fewer samples are publicly available for training. This creates a fundamental challenge: good detection performance is essential, but data scarcity negatively impacts the detection performance. Potential remedies include augmenting datasets with synthetic samples or exploring few-shot approaches (as in  (Tseng2025BIV, 29)) to enable detection of multiple sensitive regions from limited examples.
@@
DLV3+(chen2018encoder, 6) 39.3 (66.0) 65.2 0.0 70.5 22.9 58.8 0.0 0.0 67.2 17.7
YOLO-s(yolov8, 16) 41.6 (58.7) 67.8 41.3 61.4 35.4 40.1 29.1 14.0 14.5 23.2
YOLO-x(yolov8, 16) 43.4 (60.6) 69.5 41.6 62.9 39.9 44.2 25.4 18.2 19.1 22.3
SFM-B5(xie2021segformer, 30) 42.1 (75.8) 78.6 0.0 79.4 46.6 66.8 0.0 0.0 55.7 0.0
FPN-R(he2017mask, 14) 43.8 (71.5) 73.6 44.7 78.1 32.5 29.6 17.8 12.1 30.6 26.6
FPN-X(he2017mask, 14) 46.6 (74.0) 74.1 53.3 80.7 34.6 32.7 24.9 11.1 37.6 26.4
@@
MPNet(song2020mpnet, 27) 80.4 68.1 76.5 69.2 91.4 8.6 78.3 72.3 80.4
DeBERTa(he2021deberta, 15) 81.6 70.5 76.9 74.4 91.7 13.0 80.6 76.0 80.8
BERT(devlin2019bert, 9) 91.1 79.8 91.3 89.6 94.8 11.2 91.7 88.7 91.6
@@
Cascade R-CNN(cai2018cascade, 5) 46.2 23.5 85.2 49.8 18.1 37.4 50.9 58.7
RetinaNet(lin2017focal, 18) 49.9 38.7 84.1 55.0 29.2 45.1 48.8 48.9
Faster R-CNN(ren2015faster, 24) 51.1 37.5 86.1 50.8 29.6 52.9 39.8 60.9
YOLO(yolov8, 16) 58.8 38.6 89.7 51.6 52.4 58.0 51.1 70.3

### v2

@@ 6.1. PSO Detection Performance
Table 3(a) presents the performance of segmentation models, highlighting variations across object categories. SegFormer-B5 and DeepLabV3+ perform particularly well on categories including face, person, handwriting, and fingerprint, but fail in classes that have fewer samples in the training set, such as license plate, physical disability, and medical history. In contrast, YOLO-Seg, FPN-R, and FPN-X provide more balanced performance results in all classes, with FPN-X101 achieving the highest overall mIoU score. Our results indicate that all models exceed a certain threshold for face and person detection, reflecting knowledge already embedded in pre-trained weights and refined during fine-tuning. However, their performance drops significantly for classes such as physical disability, signature, and medicine, largely due to class imbalance. These categories are considered highly sensitive; therefore, fewer samples are publicly available for training. This creates a fundamental challenge: good detection performance is essential, but data scarcity negatively impacts the detection performance. Potential remedies include augmenting datasets with synthetic samples or exploring few-shot approaches (as in  (Tseng et al., 2025)) to enable detection of multiple sensitive regions from limited examples.
@@
DLV3+(Chen et al., 2018) 39.3 (66.0) 65.2 0.0 70.5 22.9 58.8 0.0 0.0 67.2 17.7
YOLO-s(Jocher and others, 2023) 41.6 (58.7) 67.8 41.3 61.4 35.4 40.1 29.1 14.0 14.5 23.2
YOLO-x(Jocher and others, 2023) 43.4 (60.6) 69.5 41.6 62.9 39.9 44.2 25.4 18.2 19.1 22.3
SFM-B5(Xie et al., 2021) 42.1 (75.8) 78.6 0.0 79.4 46.6 66.8 0.0 0.0 55.7 0.0
FPN-R(He et al., 2017) 43.8 (71.5) 73.6 44.7 78.1 32.5 29.6 17.8 12.1 30.6 26.6
FPN-X(He et al., 2017) 46.6 (74.0) 74.1 53.3 80.7 34.6 32.7 24.9 11.1 37.6 26.4
@@
MPNet(Song et al., 2020) 80.4 68.1 76.5 69.2 91.4 8.6 78.3 72.3 80.4
DeBERTa(He et al., 2021) 81.6 70.5 76.9 74.4 91.7 13.0 80.6 76.0 80.8
BERT(Devlin et al., 2019) 91.1 79.8 91.3 89.6 94.8 11.2 91.7 88.7 91.6
@@
Cascade R-CNN(Cai and Vasconcelos, 2018) 46.2 23.5 85.2 49.8 18.1 37.4 50.9 58.7
RetinaNet(Lin et al., 2017) 49.9 38.7 84.1 55.0 29.2 45.1 48.8 48.9
Faster R-CNN(Ren et al., 2017) 51.1 37.5 86.1 50.8 29.6 52.9 39.8 60.9
YOLO(Jocher and others, 2023) 58.8 38.6 89.7 51.6 52.4 58.0 51.1 70.3
@@ 6.2. PSO Protection Performance
7. Security Discussion
We discuss how the proposed design addresses confidentiality and enforces access control while considering practical deployment.
Preserving Confidentiality.
The CryptoCore module encrypts each PSO with a symmetric key tied to its sensitivity group. Each symmetric key is further protected via ABE under policies defined by the AccessPolicy Module. Raw PSOs never leave the CryptoCore module; only metadata is visible to other system components.
Access Control.
User access is determined by ABE keys encoding their attributes. Decrypting a symmetric key allows access only to PSOs permitted by the corresponding policy, while higher-sensitivity PSOs remain encrypted, ensuring strict enforcement of access rules.
Collusion Considerations.
Low-risk collusions, like users sharing decrypted PSOs, do not expose sensitive content. High-risk collusions involving the AccessPolicy or CryptoCore modules could compromise confidentiality but are out of scope of this work. However, Trusted Execution Environments could potentially be used to mitigate them.
Metadata Leakage
Although raw PSOs are encrypted, metadata may leak side-channel information. We define a leakage function $\mathcal{L}$ mapping a dataset $\mathcal{D}$ to potentially exposed information:
$\mathcal{L}(\mathcal{D})=\left\{\begin{aligned} &\text{PSO positions,class frequencies},\\ &\text{confidence distributions,sensitivity group counts},\\ \end{aligned}\right\}.$
This leakage is inherent to object-detection-based systems and unavoidable without obfuscation (Barak et al., 2001) or oblivious RAM (ORAM) (Goldreich and Ostrovsky, 1996), which are impractical in this setting. In practice, ORAM adds significant bandwidth and latency overhead due to repeated oblivious memory accesses, and efficient general-purpose obfuscation does not exist beyond theoretical constructions, making both incompatible with real-time object detection pipelines.
To formalize the confidentiality guarantees and access control guarantees, we show that even with access to $\mathcal{L}(\mathcal{D})$, no adversary can gain non-negligible advantage in recovering PSO contents without satisfying the associated access policies.
Theorem 7.1 (Confidentiality & Access Control under Metadata Leakage).
If $\mathsf{SKE}$ is IND-CPA  [^2] IND-CPA: Indistinguishability under Chosen-Plaintext Attack, a standard notion of semantic security. secure and $\mathsf{ABE}$ is IND-CPA secure and collusion-resistant, then no PPT adversary $\mathcal{A}$ that cannot satisfy policy $P_{\ell}$ and observes metadata via $\mathcal{L}$ can distinguish encryptions of two chosen $\mathsf{PSO}$s from group $\ell$ with non-negligible advantage.
Proof Sketch.
Suppose a PPT adversary $\mathcal{A}$ has a non-negligible advantage $\epsilon$ in distinguishing $\mathsf{PSO}_{0}$ and $\mathsf{PSO}_{1}$, possibly using $\mathcal{L}(\mathcal{D})$. We construct a PPT adversary $\mathcal{B}$ that breaks $\mathsf{SKE}$ or $\mathsf{ABE}$:
  • If $\mathcal{A}$ never recovers the symmetric key $\mathsf{K}_{\ell}$, distinguishing $c_{\mathsf{PSO}_{b}}\leftarrow\mathsf{SKE.Enc}(\mathsf{K}_{\ell},\mathsf{PSO}_{b})$ gives $\mathcal{B}$ an IND-CPA attack on $\mathsf{SKE}$ with advantage $\epsilon$.
  • If $\mathcal{A}$ obtains $\mathsf{K}_{\ell}$ without satisfying $P_{\ell}$, $\mathcal{B}$ uses this to distinguish $c_{P_{\ell}}$ under $\mathsf{ABE}$ IND-CPA, breaking $\mathsf{ABE}$ with advantage $\epsilon$.
  • Metadata from $\mathcal{L}(\mathcal{D})$ reveal, at most positions, class frequencies, or sensitivity counts. Since it does not expose plaintext PSOs or symmetric keys, it contributes only negligible advantage.
Thus, in all cases, $\mathcal{B}$ contradicts the assumed security of $\mathsf{SKE}$ or $\mathsf{ABE}$, and metadata leakage is negligible. Therefore, $\epsilon$ is negligible.
PSO Leakage due to Detection Misses
As shown in Section 6, ML models can still produce false negatives, i.e., fail to detect certain PSOs, even when an integrated ensemble approach is used. Achieving zero false negatives is challenging with current general-purpose segmentation and object detection models. Therefore, sensitive objects or pixels may remain unencrypted, potentially leading to privacy leakage. To mitigate this limitation, we propose to enable user control over the detection and classification module. In the context of the proposed architecture, this functionality can be incorporated into the Post-Correction Module. Such user interaction and correction mechanisms can support the refinement and further fine-tuning of the ML models, thereby improving control over privacy leakage. The integration of human feedback into the architecture, as well as the development of human validation methods to assess privacy risks due to false negatives, is left for future work. We consider this direction crucial for enhancing the practical applicability of the proposed architecture and for increasing user trust in automated access control systems.